# CityBench: Evaluating the Capabilities of Large Language Models for Urban Tasks

### v2

@@ Abstract
Recently, large language models (LLMs) with extensive general knowledge and powerful reasoning abilities have seen rapid development and widespread application. A systematic and reliable evaluation of LLMs or vision-language model (VLMs) is a crucial step in applying and developing them for various fields. There have been some early explorations about the usability of LLMs for limited urban tasks, but a systematic and scalable evaluation benchmark is still lacking. The challenge in constructing a systematic evaluation benchmark for urban research lies in the diversity of urban data, the complexity of application scenarios and the highly dynamic nature of the urban environment. In this paper, we design CityBench, an interactive simulator based evaluation platform, as the first systematic benchmark for evaluating the capabilities of LLMs for diverse tasks in urban research. First, we build CityData to integrate the diverse urban data and CitySimu to simulate fine-grained urban dynamics. Based on CityData and CitySimu, we design 8 representative urban tasks in 2 categories of perception-understanding and decision-making as the CityBench. With extensive results from 30 well-known LLMs and VLMs in 13 cities around the world, we find that advanced LLMs and VLMs can achieve competitive performance in diverse urban tasks requiring commonsense and semantic understanding abilities, e.g., understanding the human dynamics and semantic inference of urban images. Meanwhile, they fail to solve the challenging urban tasks requiring professional knowledge and high-level reasoning abilities, e.g., geospatial prediction and traffic control task. Codes are openly accessible via https://github.com/tsinghua-fib-lab/CityBench.
1 Introduction
@@
Recent years, large language models (LLMs) with extensive commonsense and reasoning capabilities have achieved excellent results in various fields (Achiam et al., 2023; Touvron et al., 2023), including programming (Hong et al., 2023), mathematics (Wei et al., 2022), visual intelligence (Liu et al., 2024c) and commonsense reasoning (Suzgun et al., 2022; Mialon et al., 2023). Furthermore, powerful LLMs enable many unimaginable research endeavors to become feasible, e.g., agent (Wang et al., 2024a), embodied intelligence (Reed et al., 2022) and AGI (Feng et al., 2024b). These researchers postulate that LLMs, by acquiring extensive world knowledge and commonsense, hold the key to unlocking promising outcomes in these challenging applications. Many works (Achiam et al., 2023; Gurnee & Tegmark, 2023; Roberts et al., 2023) have demonstrated that LLMs can be regarded as ‘world models’ of our life and they are skilled at solving a wide variety of tasks across multiple fields, while other works (Xiang et al., 2024; Yang et al., 2023; Wang et al., 2024b) indicate that LLMs lack an comprehensive understanding of the real physical world and fail to handle many real-life problems. However, these research efforts have primarily focused on the indoor environment (Puig et al., 2018), while neglecting the outdoor environment, specifically the broader urban environment (Batty et al., 2012; Zheng et al., 2014).
@@
Various works have explored the potential of LLMs in modeling urban space and solving urban tasks. For example, researchers evaluate the potential of LLMs on remote sensing understanding tasks (Kuckreja et al., 2023; Zhang & Wang, 2024) and urban visual tasks (Yan et al., 2024; Hao et al., 2024). Gurnee et al. (Gurnee & Tegmark, 2023) and Roberts et al. (Roberts et al., 2023) evaluate whether LLMs acquire the spatial knowledge of the world, such as cities and coordinates. Manvi et al. (Manvi et al., 2023; 2024) try to extract the geospatial knowledge in LLMs to conduct geospatial indicator prediction tasks (Mai et al., 2023). Besides, researchers also explore how to apply LLMs into the realistic urban applications, e.g., traffic control (Da et al., 2023; Lai et al., 2023), traffic prediction (Liu et al., 2024a), mobility prediction (Wang et al., 2023b), visual language navigation (Schumann et al., 2024) and so on. However, on the one hand, these existing works primarily focus on evaluating the static spatial knowledge of LLMs without considering the environment dynamics and interactivity. On the other hand, most of them only focus on one type of task and one modality of data in the urban space, using small dataset that are not scalable globally. Although there are some existing simulators for urban space such as game simulators (Interactive, 2023) and traffic simulators (Lopez et al., 2018), they cannot be directly applied to support the evaluation and significant amount of adaptation work is required. None of them can support the systematic evaluation of LLMs’ capabilities for diverse tasks in urban research, ranging from understanding and reasoning to decision-making tasks.
@@
Figure 1: The framework of CityBench, which consists of a data collector CityData, an activity simulator CitySimu and 8 diverse urban tasks with different modalities. The evaluation data in the benchmark is collected from 13 cities around the world.
@@
  • Extensive experiments on CityBench with 30 well-known open source and commercial LLMs and VLMs demonstrate the effectiveness of CityBench as evaluation benchmark and also discuss the potential and limitation of applying LLMs and VLMs in urban tasks, ranging from understanding and reasoning to decision-making task. Our results shed light for the further application and development of LLMs and VLMs in the future.
Figure 2: The pipeline of building benchmark, including data collection stage, data integration stage, evaluation generation stage and quality control stage.
2 Related Work
LLM Evaluation for Urban Research Researchers from various urban related fields have conducted extensive evaluations of LLM in urban space from different aspects. Zhang et al (Zhang & Wang, 2024) and Kuckreja et al. (Kuckreja et al., 2023) evaluate the performance of multi-modal LLMs on several remote sensing related tasks. Yang et al. (Yang et al., 2024) propose V-IRL benchmark to evaluate the performance of multi-modal LLMs on street view image related tasks including localization and recognition tasks. Mai et al. (Mai et al., 2023) and Manvi et al. (Manvi et al., 2023) use LLMs to predict social indicators like population and education level. Gurnee et al. (Gurnee & Tegmark, 2023) and Bhandari et al. (Bhandari et al., 2023) try to testify whether LLMs know the coordinates of geospatial entity. Mooney et al. (Mooney et al., 2023) and Deng et al. (Deng et al., 2023) use GIS exams to understand the geospatial skills of LLMs. Roberts et al. (Roberts et al., 2023) design GPT4GEO to evaluate the geospatial capabilities of LLM with limited case studies. Different from these works, we first introduce the interactive simulator based systematic evaluation system for LLMs and VLMs, which covers various data modalities, diverse urban task types and differentiated data from 13 cities around the world.
Interactive Decision-making and Urban Simulator Beyond the above static evaluation, researchers also evaluate the capacity of LLMs in the interactive decision making tasks with customized simulators, e.g., web agent (Liu et al., 2023; Zeng et al., 2023) with web environment and embodied intelligence (Yang et al., 2023) with virtual home (Puig et al., 2018). In the urban domain, Schumann et al. (Schumann et al., 2024) apply LLM to do the visual language navigation task in Touchdown (Chen et al., 2019) and Lai (Lai et al., 2023) apply LLMs as the traffic light controller in CityFlow (Zhang et al., 2019) to manage the road traffic. Besides, Yang et al. (Yang et al., 2024) design V-IRL as the environment of street view image related tasks and propose a global scale virtual intelligence benchmark. These works only evaluate the potential of LLMs in single urban decision-making task and most of their results rely on small-scale datasets in limited regions. Different from them, our work builds on an efficient urban simulator with global scale and supports 4 representative urban decision-making tasks with different modality in one benchmark, including urban exploration, outdoor navigation, mobility prediction and traffic control task.
3 Methods
@@
3.1 CityData
To present a complete picture of the city’s geospatial structure, semantic features, and human activities, CityData integrates the following globally available data from multiple sources.
Geospatial Data Geospatial data, represented by maps, is the most fundamental data for describing the urban structure including road networks, points of interest (POIs), areas of interest (AOIs), etc. OpenStreetMap (OSM) ***https://openstreetmap.org/ is most widely used open source map data. However, the raw data provided by OSM cannot support the simulation of urban dynamics directly because the relationship between different elements is incomplete such as the connection between buildings and roads. Therefore, we provide a globally available rule-based map building tool within CityData that reconstructs lanes, lane topology, and building-lane connections based on the raw OSM data. The reconstructed map is used as the geospatial base and simulation input in CitySimu.
Urban Visual Data Street view data and satellite images are two types of globally available urban data that contains rich semantic information, which represents the visual of human. Therefore, CityData also integrates the two types of data, the former obtained via Google Maps API and Baidu Maps API, and the latter using the Esri World Imagery as data source. In the open-source CityData python package, street view data is accessed through spatial location and facing direction, and satellite images are acquired through spatial ranges.
Human Activities Data We use the global Foursquare-checkin (Yang et al., 2016) data and a synthetic global origin-destination data (OD data) (loooffeeeey, 2024) as the proxy of human activities to enable the fine-grained human movement simulation. The Foursquare-checkin dataset (Yang et al., 2016) is a long-term user check-in dataset collected from Foursquare https://foursquare.com/ in approximately 400 cities worldwide. It has been widely used in the community over the past ten years (Chen et al., 2024). Origin-destination data is generated by a diffusion model with population from Worldpophttps://www.worldpop.org/ and satellite image from Esri World Imagery as input. While all the user information are anonymized, we follow the license from Foursquare-checkin (Yang et al., 2016) to protect the public privacy.
Figure 3: The simulation framework of CitySimu, including base environment APIs, interactive objects, simulation APIs and language APIs. Besides, supported task examples also present the relation between simulation APIs and evaluation tasks.
3.2 CitySimu
@@
Traffic Simulation In the former two simulations, we only simulate the individual actions without the interaction with others. Here, we introduce microscopic traffic simulation to model the interaction behaviors between vehicles and provide a traffic control environment. The simulator takes the geospatial data reconstructed from OSM within CityData and the travel demand described by the synthetic global OD data as inputs. It simulates the vehicle behaviors through realistic driving simulation models including the intelligent driver model (IDM) (Treiber et al., 2000) as the car-following model and the randomized MOBIL model (Kesting et al., 2007; Feng et al., 2021) as the lane-change model to obtain the dynamics of all vehicles in the city at each second. Through the sensing APIs, LLMs can obtain data about urban dynamics such as junction queue length, vehicle speed, and road average speed. Through the control APIs, LLMs can intervene in the city’s operation, such as modifying traffic signal lights, modifying the speed limit of the road, etc.
3.3 CityBench
Based on CityData and CitySimu, we design a multi-modal urban evaluation benchmark CityBench to evaluate the capability of LLMs and VLMs. In the following section, we first summarize the whole pipeline and then give introduction to each task. Due to the limits of space, detailed templates and prompts for each task can refer to the appendix.
3.3.1 Pipeline
@@
In the evaluation generation stage, we use template based methods and LLMs/VLMs based methods to generate the evaluation questions. For example, for the image geolocalization task, the groundth location is already known when collecting, thus we directly design template based question to convert the image geolocalization task into question answer pair. As for the outdoor navigation task, we employ VLM to act as human annotation experts to annotate the data to generate the navigation instruction with additional inputs. In CityBench, the instructions for urban exploration task and the outdoor navigation task are generated by LLM assisted methods. Instructions for other tasks are generated from template based methods.
Due to the hand-craft designs and potential issues of LLMs, we apply a quality control stage to filter and rewrite the generated questions to obtain a high quality evaluation questions. For questions generated from template based methods, we use LLM as data quality expert to filter the low-quality data and use LLM as data rewritter to rewrite the questions with diverse formats and expressions. For questions generated from the LLMs/VLMs based methods, we use LLM/VLM as the agent with additional information to execute the task to verify the quality of generated instructions. If the generated questions are filtered too much, we will return to the evaluation generation stage to generate new questions again. Finally, the authors of this paper also participate in the quality control stage to filter and rewrite the generated data to ensure the quality of whole benchmark.
@@
Figure 4: 8 tasks in CityBench with their metrics.
3.3.2 Perception and Understanding Task
Image Geolocalization Image geolocalization task is to predict the precise location of image based on its context. Street view image is regarded as the recording of urban appearance and play an important role in understanding the urban environment and dynamics (Fan et al., 2023). Thus, we query VLMs with street view image and require them to directly generate the location of image. A good VLM should recognize the important objects from the street image and mapping them into the potential locations. Following  (Haas et al., 2023), we define two subtasks for this task: city name inference and precise latitude and longitude coordinates inference.
@@ 3.3.2 Perception and Understanding Task
Infrastructure Inference Besides, we also introduce the infrastructure inference task which means to recognize the urban infrastructures from the satellite images. The groundtruth of this task is extracted from the OSM by matching predefined infrastructure key words within a fixed spatial range. Given the satellite image and a list of all kinds of infrastructures, VLM is required to generate the infrastructure names appeared in the image.
GeoQA for City Elements Beyond understanding the urban space from the visual perspective, we introduce geographic question answer(GeoQA) (Mai et al., 2021) to test whether LLMs comprehends the fundamental spatial elements (Lynch, 1964) in a city from the concept view, such as road, landmarks and boundary. Following (Lynch, 1964; Feng et al., 2024a), we classify the spatial elements into six groups and design problems for each group.
3.3.3 Planning and Decision Making Task
@@
Mobility Prediction As one of the fundamental task for understanding the human behaviors and urban dynamics, mobility prediction task is to predict the next location of user in the next time window with given the past mobility trajectory. Here we use the the global Foursquare checkin data to support the mobility prediction in the simulator. We follow (Wang et al., 2023b) to conduct the mobility prediction task via LLMs.
@@
4 Benchmark and Experiments
Table 1: Multi-source data for 13 cities around the world. The spatial coverage of these cities can be found in the appendix.
World Cities Visual Data GeoSpatial Data Human Activity Data

### v3

@@ Abstract.
As large language models (LLMs) continue to advance and gain widespread use, establishing systematic and reliable evaluation methodologies for LLMs and vision-language models (VLMs) has become essential to ensure their real-world effectiveness and reliability. There have been some early explorations about the usability of LLMs for limited urban tasks, but a systematic and scalable evaluation benchmark is still lacking. The challenge in constructing a systematic evaluation benchmark for urban research lies in the diversity of urban data, the complexity of application scenarios and the highly dynamic nature of the urban environment. In this paper, we design CityBench, an interactive simulator based evaluation platform, as the first systematic benchmark for evaluating the capabilities of LLMs for diverse tasks in urban research. First, we build CityData to integrate the diverse urban data and CitySimu to simulate fine-grained urban dynamics. Based on CityData and CitySimu, we design 8 representative urban tasks in 2 categories of perception-understanding and decision-making as the CityBench. With extensive results from 30 well-known LLMs and VLMs in 13 cities around the world, we find that advanced LLMs and VLMs can achieve competitive performance in diverse urban tasks requiring commonsense and semantic understanding abilities, e.g., understanding the human dynamics and semantic inference of urban images. Meanwhile, they fail to solve the challenging urban tasks requiring professional knowledge and high-level numerical abilities, e.g., geospatial prediction and traffic control task. These findings provide critical insights for the effective utilization and further development of LLMs to advance urban-related tasks and research in the future. The associated data and code are publicly available at https://github.com/tsinghua-fib-lab/CityBench.
copyright: noneconference: Submission #15; KDD Datasets and Benchmarks; 202511footnotetext: These authors contributed equally.22footnotetext: These authors contributed equally.33footnotetext: Corresponding author, email: liyong07@tsinghua.edu.cn
1. Introduction
@@
Recent years, large language models (LLMs) with extensive commonsense and reasoning capabilities have achieved excellent results in various fields (Achiam et al., 2023; Touvron et al., 2023), including programming (Hong et al., 2023), mathematics (Wei et al., 2022), visual intelligence (Liu et al., 2024b) and commonsense reasoning (Suzgun et al., 2022; Mialon et al., 2023). Furthermore, powerful LLMs enable many unimaginable research endeavors to become feasible, e.g., agent (Wang et al., 2024a) and embodied intelligence (Reed et al., 2022; Zha et al., 2025). These researchers postulate that LLMs, by acquiring extensive world knowledge and commonsense, hold the key to unlocking promising outcomes in these challenging applications. Many works (Achiam et al., 2023; Gurnee and Tegmark, 2023) have demonstrated that LLMs can be regarded as ‘world models’ of our life and they are skilled at solving a wide variety of tasks across multiple fields, while other works (Xiang et al., 2024; Yang et al., 2023; Wang et al., 2024b) indicate that LLMs lack an comprehensive understanding of the real physical world and fail to handle many real-life problems. However, these research efforts have primarily focused on the indoor environment (Puig et al., 2018), while neglecting the outdoor environment, specifically the broader urban environment (Batty et al., 2012; Zheng et al., 2014).
@@
Various works have explored the potential of LLMs in modeling urban space and solving urban tasks. For example, researchers evaluate the potential of LLMs on remote sensing understanding tasks (Kuckreja et al., 2023) and urban visual tasks (Yan et al., 2024). Gurnee et al. (Gurnee and Tegmark, 2023) evaluate whether LLMs acquire the spatial knowledge of the world, such as cities and coordinates. Manvi et al. (Manvi et al., 2023, 2024) try to extract the geospatial knowledge in LLMs to conduct geospatial indicator prediction tasks (Mai et al., 2023). Besides, researchers also explore how to apply LLMs into the realistic urban applications, e.g., traffic control (Lai et al., 2023), mobility prediction (Wang et al., 2023a; Feng et al., 2025a), behavior modeling (Gong et al., 2024), visual language navigation (Schumann et al., 2024) and so on. However, on the one hand, these existing works primarily focus on evaluating the static spatial knowledge of LLMs without considering the environment dynamics and interactivity. On the other hand, most of them only focus on one type of task and one modality of data in the urban space, using small dataset that are not scalable globally. Although there are some existing simulators for urban space such as game simulators (Interactive, 2023) and traffic simulators (Lopez et al., 2018), they cannot be directly applied to support the evaluation and significant amount of adaptation work is required. None of them can support the systematic evaluation of LLMs’ capabilities for diverse tasks in urban research, ranging from understanding and reasoning to decision-making tasks.
@@
Figure 1. The framework of CityBench, which consists of a data collector CityData, an activity simulator CitySimu and 8 diverse urban tasks with different modalities. The evaluation data in the benchmark is collected from 13 cities around the world.
@@
  • Extensive experiments on CityBench with 30 well-known open source and proprietary LLMs and VLMs demonstrate the effectiveness of CityBench as evaluation benchmark and also discuss the potential and limitation of applying LLMs and VLMs in urban tasks, ranging from understanding and reasoning to decision-making task.
Figure 2. The pipeline of building benchmark, including data collection stage, data integration stage, evaluation generation stage and quality control stage.
2. Methods
@@
2.1. CityData
In the section, we introduce the multi-source urban dataset CityData collected to support multi-modal urban tasks. To present a complete picture of the city’s geospatial structure, semantic features, and human activities, CityData integrates the following globally available data from multiple sources (Liu et al., 2023a). The python package for data collection and access is open-source [^1] https://github.com/tsinghua-fib-lab/pycitydata.
Geospatial Data Geospatial data, represented by maps, is the most fundamental data for describing the urban structure including road networks, points of interest (POIs), areas of interest (AOIs), etc. OpenStreetMap (OSM)  [^2] https://www.openstreetmap.org/ is most widely used open source map data. However, the raw data provided by OSM cannot support the simulation of urban dynamics directly because the relationship between different elements is incomplete such as the connection between buildings and roads. Therefore, we provide a globally available rule-based map building tool   [^3] https://github.com/tsinghua-fib-lab/mosstool within CityData that reconstructs lanes, lane topology, and building-lane connections based on the raw OSM data. The reconstructed map is used as the geospatial base and simulation input in CitySimu.
Urban Visual Data Street view data and satellite images are two types of globally available urban data that contains rich semantic information, which represents the visual of human. Therefore, CityData also integrates the two types of data, the former obtained via Google Maps API and Baidu Maps API, and the latter using the Esri World Imagery as data source. In CityData, street view data is accessed through spatial location and facing direction, and satellite images are acquired through spatial ranges.
Human Activities Data We use the global Foursquare-checkin (Yang et al., 2016) data and a synthetic global origin-destination data (OD data)  [^4] https://github.com/tsinghua-fib-lab/generate-od-pubtools as the proxy of human activities to enable the fine-grained human movement simulation. The Foursquare-checkin dataset (Yang et al., 2016) is a long-term user check-in dataset collected from Foursquare  [^5] https://foursquare.com/ in approximately 400 cities worldwide. It has been widely used in the community over the past ten years (Chen et al., 2024). Origin-destination data is generated by a diffusion model with population from Worldpop  [^6] https://www.worldpop.org/ and satellite image from Esri World Imagery as input. While all the user information are anonymized, we follow the license from Foursquare-checkin (Yang et al., 2016) to protect the public privacy.
Figure 3. The simulation framework of CitySimu, including base environment APIs, interactive objects, simulation APIs and language APIs. Besides, supported task examples also present the relation between simulation APIs and evaluation tasks.
2.2. CitySimu
@@
Traffic Simulation In the former two simulations, we only simulate the individual actions without the interaction with others. Here, we introduce microscopic traffic simulation to model the interaction behaviors between vehicles and provide a traffic control environment. The simulator takes the geospatial data reconstructed from OSM within CityData and the travel demand described by the synthetic global OD data as inputs. It simulates the vehicle behaviors through realistic driving simulation models including the intelligent driver model (IDM) (Treiber et al., 2000) as the car-following model and the randomized MOBIL model (Kesting et al., 2007; Feng et al., 2021) as the lane-change model to obtain the dynamics of all vehicles in the city at each second. The simulator also provides a series of sensing and control APIs. Through the sensing APIs, LLMs can obtain data about urban dynamics such as junction queue length, vehicle speed, and road average speed. Through the control APIs, LLMs can intervene in the city’s operation, such as modifying traffic signal lights, modifying the speed limit of the road, etc.
2.3. CityBench
Based on CityData and CitySimu, we design a multi-modal urban evaluation benchmark CityBench to evaluate the capability of LLMs and VLMs. In the following section, we first summarize the whole pipeline and then give introduction to each task.
2.3.1. Pipeline
@@
In the evaluation generation stage, we use template based methods and LLMs/VLMs based methods to generate the evaluation questions. For example, for the image geolocalization task, the groundth location is already known when collecting, thus we directly design template based question to convert the image geolocalization task into question answer pair. As for the outdoor navigation task, we employ VLM to act as human annotation experts to annotate the data to generate the navigation instruction with additional inputs. In CityBench, instructions for urban exploration task and outdoor navigation task are generated by LLM assisted methods. Instructions for other tasks are generated from template based methods.
Due to the hand-craft designs and potential issues of LLMs, we apply a quality control stage to filter and rewrite the generated questions to obtain a high quality evaluation questions. For questions generated from template based methods, we use LLM as data quality expert to filter the low-quality data and use LLM as data rewritter to rewrite the questions with diverse formats and expressions. For questions generated from the LLMs/VLMs based methods, we use LLM/VLM as the agent with additional information to execute the task to verify the quality of generated instructions. If the generated questions are filtered too much, we will return to the evaluation generation stage to generate new questions again. Finally, authors of this paper also participate in the quality control stage to filter and rewrite the generated data to ensure the quality of whole benchmark.
@@
Figure 4. 8 tasks in CityBench with their metrics.
2.3.2. Perception and Understanding Task
The first task is the street view image geolocalization task from the urban visual intelligence (Fan et al., 2023). Following are social indicator prediction and infrastructure inference tasks from remote sensing field. Finally, we adapt GeoQA (Mai et al., 2021; Feng et al., 2024) task into urban environment.
Image Geolocalization Image geolocalization task is to predict the precise location of image based on its context. Street view image is regarded as the recording of urban appearance and play an important role in understanding the urban environment and dynamics (Fan et al., 2023). Thus, we query VLMs with street view image and require them to directly generate the location of image. A good VLM should recognize the important objects from the street image and mapping them into the potential locations. Following  (Haas et al., 2023), we define two subtasks for this task: city name inference and precise latitude and longitude inference.
@@ 2.3.2. Perception and Understanding Task
Infrastructure Inference Besides, we also introduce the infrastructure inference task which means to recognize the urban infrastructures from the satellite images. This task require the ability of scene understand and object segmentation of urban environment. The groundtruth of this task is extracted from the OSM by matching predefined infrastructure key words within a fixed spatial range. Given the satellite image and a list of all kinds of infrastructures, VLM is required to generate the infrastructure names appeared in the image. Here, we pay attention to the following infrastructures: Airport, Harbor, Stadium, Bridge, Roundabout and Train Station.
GeoQA for City Elements Beyond understanding the urban space from the visual perspective, we introduce geographic question answer(GeoQA) (Mai et al., 2021) to test whether LLMs comprehends the fundamental elements (Lynch, 1964) in a city from the concept view, such as road and landmarks. For example, we directly ask LLM about the relation between different roads in a city. Following (Lynch, 1964; Feng et al., 2024), we classify the spatial elements into six groups and design problems for each group. These six groups are node, path, landmark, boundary, districts and others.
2.3.3. Planning and Decision Making Task
@@
Mobility Prediction As one of the fundamental task for understanding the human behaviors and urban dynamics, mobility prediction task is to predict the next location of user in the next time window with given the past mobility trajectory. Here we use the the global Foursquare checkin data to support the mobility prediction in the simulator. We follow (Wang et al., 2023a) to conduct the mobility prediction task via LLMs.
@@
3. Benchmark and Experiments
Table 1. Detailed statistics of multi-source data for 13 global cities utilized in CityBench.
Cities Visual Data GeoSpatial Data Human Activity Data